# ReCode: Unify Plan and Action for Universal Granularity Control

### v2

@@ 1 Introduction
Human decision-making in the real world naturally operates across varying granularities, seamlessly integrating high-level planning with fine-grained actions. Consider the simple act of preparing breakfast: one effortlessly shifts from deciding a high-level plan like “making bacon and eggs” to executing detailed, motor-level actions such as “cracking an egg”. Cognitive science attributes this fluid adaptability to the human brain’s integrated approach to control, which allows for decisions to be managed across different granularities. This integration is reflected in the brain’s shared representations for related cognitive processes (Prinz, 1997) and its hierarchical structure (Koechlin et al., 2003; Badre & D’esposito, 2009). Thus, humans naturally master this dynamic control of decision granularity, an essential capability for adaptive, intelligent behavior.
Achieving this adaptive intelligence is a primary goal for LLM-based agents (Liu et al., 2025), yet current frameworks fall short because they operate at fixed granularities. The core issue is that existing approaches explicitly separate planning from action into distinct decision-making processes. For example, ReAct (Yao et al., 2023) agent as shown in Figure 1(a), alternates strictly between reasoning and primitive actions step-by-step, limiting it to fine-grained decision-making without strategic foresight. On the other hand, agent with planner module (Figure 1(b)) separates high-level planning from low-level action using predefined structures. Such rigid boundaries impede agents from dynamically adjusting their decision granularity in response to evolving task complexities, ultimately resulting in brittle performance in complex, real-world environments.
@@
Extensive experiments conducted across diverse, complex decision-making benchmarks validate the effectiveness of the ReCode paradigm. ReCode consistently surpasses strong paradigm baselines such as ReAct (Yao et al., 2023) and CodeAct (Wang et al., 2024b), delivering substantial improvements in overall task performance and problem-solving efficiency. Furthermore, our experiments reveal that the hierarchical, multi-granular data generated by ReCode plays a critical role in training flexible decision-making capabilities. These empirical results demonstrate ReCode’s potential as a transformative approach to building more adaptive, capable, and generalizable LLM-based agents.
@@ 2 Related Work
The powerful reasoning and language capabilities of Large Language Models (LLMs) have inspired significant research effort to create agents that can interact with complex environments and accomplish specific tasks (Ahn et al., 2022; Huang et al., 2023; Yao et al., 2023). These efforts have largely fallen into two dominant paradigms: agents following the ReAct paradigm (Yao et al., 2023) and agents with explicit planners. However, both approaches, in their current forms, struggle with the fundamental limitation of a rigid separation between high-level planning and low-level action, which prevents flexible decision granularity control.
@@
Among these, paradigms like ReAct (Yao et al., 2023) are foundational, interleaving reasoning and primitive actions in a step-by-step loop. Code has also emerged as an expressive medium for actions, with paradigms like CodeAct (Wang et al., 2024b) leveraging the extensive code pre-training of LLMs. These foundational paradigms have spawned numerous agent systems (Yang et al., 2023a; Wu et al., 2024; Zhang et al., 2025; Liang et al., 2025; Hu et al., 2025). Much of the subsequent research has focused on optimizing this step-by-step loop, such as through synthetic data generation methods (Zelikman et al., 2022; Yang et al., 2024b; Wang et al., 2024a) and novel training algorithms (Song et al., 2024; Qiao et al., 2024; Du et al., 2024; Lee et al., 2024; Xia et al., 2025; Xiong et al., 2025; Cao et al., 2025). While these optimization efforts improve performance, they do not address the paradigm’s fundamental lack of integrated, high-level planning, which limits its foresight in complex, long-horizon tasks.
@@
Planning emerged as a distinct research thrust, specifically to address this lack of foresight in ReAct Agent paradigm. Early approaches often adopted a “plan-and-execute” strategy, where a complete natural language plan is generated upfront before any actions are taken (Wang et al., 2023; Yang et al., 2023b; Kagaya et al., 2024; Sun et al., 2024a). However, these static plans are often brittle and difficult to adapt to dynamic environments. To address this, more sophisticated methods emerged, such as hierarchical planning frameworks (Paranjape et al., 2023; Sun et al., 2024b) and dynamic re-planning methods (Sun et al., 2023; Prasad et al., 2024), which continuously modify their plans based on environmental feedback. Nonetheless, these traditional planning methods still share the fundamental flaw of maintaining a rigid separation between the high-level plan and the low-level action execution.
While some recent approaches have begun to bridge this gap by explicitly integrating recursion or code-based processes (Liu et al., 2024; Schroeder et al., 2025; Zhang & Khattab, 2025), they still do not achieve a full unification and adaptive granularity control. It is highlighted that the core limitation across all existing paradigms remains this rigid separation of plan and action. In contrast, our work is built on the key insight that plan and action are not distinct processes, but rather represent decisions at different levels of granularity. This forms the foundation for our paradigm, which is unifying plan and action in a single representation.
@@ 4.1 Experimental Setup
We conduct experiments across three text-simulated environments that represent diverse decision-making challenges for LLM-based agents. ALFWorld (Shridhar et al., 2021) provides an embodied household environment where agents navigate and manipulate objects to complete daily tasks. WebShop (Yao et al., 2022) simulates an online shopping scenario requiring agents to search, compare, and purchase products based on specific requirements. ScienceWorld (Wang et al., 2022) presents a scientific laboratory setting where agents conduct experiments and manipulate scientific instruments. All three environments are partially observable Markov decision processes where agents cannot fully infer the environment state from observations alone. For more detailed descriptions of the environments, please refer to Appendix A.1.
@@
For inference experiments, we compare against several mainstream paradigms: ReAct (Yao et al., 2023), which alternates between reasoning and action for iterative environment interaction; CodeAct (Wang et al., 2024b), extending ReAct’s action space from natural language to Python code. And some of the work focused on improving LLM-based agent planning: AdaPlanner (Sun et al., 2023), which pre-writes action sequences in Python and iteratively modifies them based on environmental feedback; and ADaPT (Prasad et al., 2024), which decomposes tasks on-demand using ReAct as the sub-task executor.
@@
For training experiments, which evaluate how well a paradigm’s data structure lends itself to learning, we conduct supervised fine-tuning (SFT) (Wei et al., 2022) experiments on ReCode, ReAct, and CodeAct to directly compare their learning efficiency. Also, we compare some advanced training methods on the ReAct-style agent, including ETO (Song et al., 2024), which learns from expert trajectories via Direct Preference Optimization (Rafailov et al., 2023); and WKM (Qiao et al., 2024), introducing parameterized world models for training on expert and synthetic trajectories. Due to differences in models, training setups, or environment configurations, we could not directly implement these latter works, but list them for reference.
@@
For inference experiments, all methods use GPT-4o mini (Hurst et al., 2024) to ensure fair comparison. We adapt few-shot prompts from prior work, making minimal modifications to fit each method’s format while keeping the number of examples consistent. The few-shot examples used for ReCode are provided in Appendix A.3.
For training experiments, we use Qwen2.5-7B-Instruct (Yang et al., 2024a) as the base model. We construct the training datasets through the following process: First, for each training instance, we use a powerful teacher model, DeepSeek-V3.1 (DeepSeek-AI, 2024), to generate trajectories using ReAct, CodeAct, and ReCode respectively. Second, we filter these trajectories by keeping only the top 40% based on final reward. Third, from each retained trajectory, we extract input-output pairs as SFT training data, where the input excludes few-shot examples to focus on the model’s learning of the paradigm itself.
Table 1 presents detailed statistics of the resulting training datasets. As shown, different paradigms produce vastly different amounts of training data from the same pool of source instances. ReCode generates significantly fewer but more compact data pairs (e.g., 6,385 pairs with 4.96M tokens on ALFWorld) compared to ReAct (27,607 pairs with 18.96M tokens). All models are trained using full-parameter fine-tuning on 8 NVIDIA A800 80GB GPUs with the open-source verl framework (Sheng et al., 2024).
@@ 4.2 Inference Results
As shown in Table 2, ReCode achieved significant performance improvements across all three environments, with an average score of 60.8, surpassing the best baseline method by 10.5 (relative 20.9%). Notably, compared to the ReAct and CodeAct paradigms, ReCode improved average performance by 13.4 and 16.9, respectively. The results demonstrate ReCode’s exceptional capabilities across diverse task domains. In the ALFWorld environment, our method exhibits particularly strong performance, achieving a high score of 83.57 on seen tasks and an outstanding 96.27 on unseen tasks, substantially outperforming all baseline methods. On WebShop, ReCode maintains competitive performance with a score of 39.97, representing a 21.9% improvement over ADaPT. In the ScienceWorld environment, while ReCode shows comparable performance to ReAct on seen data and notable improvement on unseen tasks. These consistent improvements across varied environments underscore the effectiveness of our approach in enhancing both task-specific performance and cross-domain adaptability. We provide detailed case studies in Section 4.4, showcasing representative trajectories from ALFWorld to illustrate ReCode’s decision-making process.
@@
To illustrate that the ReCode paradigm is applicable to most models, we also experiment with Gemini 2.5 Flash (Comanici et al., 2025) and DeepSeek-V3.1, with results shown in Table 2. The cross-model evaluation demonstrates ReCode’s robustness and generalizability beyond a single LLM. On Gemini 2.5 Flash, ReCode achieved an average performance of 66.2, maintaining consistent improvements over baseline methods. Similarly, when applied to DeepSeek-V3.1, it obtained a 69.2 average score, further validating the paradigm’s effectiveness. These results indicate that ReCode’s performance gains are not dependent on specific model, but rather stem from the fundamental improvements in the reasoning and code generation process.
@@
As shown in Table 3, ReCode is remarkably cost-efficient. On average, a ReCode trajectory costs 78.9% less than ReAct and 84.4% less than CodeAct. This dramatic improvement stems from ReCode’s structured exploration. Instead of the extensive, token-intensive, step-by-step reasoning cycles characteristic of reactive methods, ReCode’s hierarchical decomposition allows the agent to make more abstract and potent decisions. This leads to shorter, more direct reasoning paths, significantly reducing the number of API calls and total tokens required to solve a task while achieving superior performance.
@@ 4.3 Training Results
This section demonstrates the effectiveness of ReCode’s hierarchical data structure for agent training. We present comprehensive results across three environments and analyze the relationship between data and performance.
@@
Table 5: Training data statistics and corresponding ScienceWorld performance for ReAct and ReCode. The table lists the number of data pairs and the resulting average rewards for each method at different filtering percentiles ().
4.4 Case Study
To provide a concrete illustration of ReCode’s decision-making process, we analyze a representative trajectory from the ALFWorld environment for the task “put two alarmclock in dresser”. The complete execution trace and a diagram of the code generation flow are available in Figure 5 of the Appendix A.4.
Initially, the agent formulates hybrid decisions that integrate abstract sub-task placeholder functions with concrete, primitive actions. These initial decisions, shown in the following code block, include functions like find_and_take for complex procedures that require further decomposition, and known, simple steps like run('go to dresser 1') as primitive actions.
The agent then executes this hybrid code block sequentially. Primitive actions like run(...) are executed directly by the environment executor. While, placeholder functions such as find_and_take trigger an on-demand expansion, where the agent generates the necessary code to complete that specific sub-task just in time. This hybrid strategy allows the agent to maintain a high-level, strategic view of the task while committing to concrete actions only when the steps are simple and certain. It combines the robustness of hierarchical planning with the efficiency of direct execution, providing a flexible and powerful problem-solving framework that can adapt its level of abstraction as needed.
@@ 6 Limitations and Future Work
While ReCode demonstrates significant potential, its effectiveness depends critically on the base model’s reasoning capabilities and few-shot example quality. The paradigm requires models to generate structured code within a specific framework, and any deficiency in planning ability compromises subsequent decision quality. As observed in our experiments, ReCode’s performance can be inconsistent, largely due to models’ difficulty adhering to the required code generation format or fully grasping the framework’s underlying logic.
These limitations point to clear research directions centered on agent learning. A primary focus should be enhancing models’ intrinsic ability to understand and operate within the ReCode framework through specialized pre-training objectives or comprehensive fine-tuning strategies. Reinforcement learning approaches could reward agents for creating efficient hierarchical plans, enabling them to optimize their own expansion processes. Additionally, exploring methods to make the framework more robust to generation errors and investigating automatic curriculum learning for progressive skill acquisition represent promising avenues. These directions would represent significant steps toward more autonomous and capable agent learning systems.

### v3

@@ 1 Introduction
Human decision-making in the real world naturally operates across varying granularities, seamlessly integrating high-level planning with fine-grained actions. Consider the simple act of preparing breakfast: one effortlessly shifts from deciding a high-level plan like “making bacon and eggs” to executing detailed, motor-level actions such as “cracking an egg”. Cognitive science attributes this fluid adaptability to the human brain’s integrated approach to control, which allows for decisions to be managed across different granularities. This integration is reflected in the brain’s shared representations for related cognitive processes (prinz1997perception) and its hierarchical structure (koechlin2003architecture; badre2009rostro). Thus, humans naturally master this dynamic control of decision granularity, an essential capability for adaptive, intelligent behavior.
Achieving this adaptive intelligence is a primary goal for LLM-based agents (liu2025advances), yet current frameworks fall short because they operate at fixed granularities. The core issue is that existing approaches explicitly separate planning from action into distinct decision-making processes. For example, ReAct (yao2023react) agent as shown in Figure 1(a), alternates strictly between reasoning and primitive actions step-by-step, limiting it to fine-grained decision-making without strategic foresight. On the other hand, agent with planner module (Figure 1(b)) separates high-level planning from low-level action using predefined structures. Such rigid boundaries impede agents from dynamically adjusting their decision granularity in response to evolving task complexities, ultimately resulting in brittle performance in complex, real-world environments.
@@
Extensive experiments conducted across diverse, complex decision-making benchmarks validate the effectiveness of the ReCode paradigm. ReCode consistently surpasses strong paradigm baselines such as ReAct (yao2023react) and CodeAct (wang2024executable), delivering substantial improvements in overall task performance and problem-solving efficiency. Furthermore, our experiments reveal that the hierarchical, multi-granular data generated by ReCode plays a critical role in training flexible decision-making capabilities. These empirical results demonstrate ReCode’s potential as a transformative approach to building more adaptive, capable, and generalizable LLM-based agents.
@@ 2 Related Work
The powerful reasoning and language capabilities of Large Language Models (LLMs) have inspired significant research effort to create agents that can interact with complex environments and accomplish specific tasks (ahn2022can; huang2023inner; yao2023react). These efforts have largely fallen into two dominant paradigms: agents following the ReAct paradigm (yao2023react) and agents with explicit planners. However, both approaches, in their current forms, struggle with the fundamental limitation of a rigid separation between high-level planning and low-level action, which prevents flexible decision granularity control.
@@
Among these, paradigms like ReAct (yao2023react) are foundational, interleaving reasoning and primitive actions in a step-by-step loop. Code has also emerged as an expressive medium for actions, with paradigms like CodeAct (wang2024executable) leveraging the extensive code pre-training of LLMs. These foundational paradigms have spawned numerous agent systems (yang2023auto; wu2024autogen; zhang2025aflow; openmanus2025; hu2025owl). Much of the subsequent research has focused on optimizing this step-by-step loop, such as through synthetic data generation methods (zelikman2022star; yang2024react; wang2024learning) and novel training algorithms (song2024trial; qiao2024agent; du2024improving; lee2024rlaif; xia2025sand; xiong2025mpo; cao2025pgpo). While these optimization efforts improve performance, they do not address the paradigm’s fundamental lack of integrated, high-level planning, which limits its foresight in complex, long-horizon tasks.
@@
Planning emerged as a distinct research thrust, specifically to address this lack of foresight in ReAct Agent paradigm. Early approaches often adopted a “plan-and-execute” strategy, where a complete natural language plan is generated upfront before any actions are taken (wang2023plan; yang2023intercode; kagaya2024rap; sun2024corex). However, these static plans are often brittle and difficult to adapt to dynamic environments. To address this, more sophisticated methods emerged, such as hierarchical planning frameworks (paranjape2023art; sun2024pearl) and dynamic re-planning methods (sun2023adaplanner; prasad2024adapt), which continuously modify their plans based on environmental feedback. Nonetheless, these traditional planning methods still share the fundamental flaw of maintaining a rigid separation between the high-level plan and the low-level action execution.
LLM-based Agent with Recursive Structure
While some recent approaches have begun to bridge this gap by explicitly integrating recursion or code-based processes, they still do not achieve a full unification and adaptive granularity control. Code-as-Policies (liang2023code) and follow-up methods (wang2023demo2code; mu2024robocodex; ahn2025towards) use complete executable programs to represent the agent policy, and recursively expand high-level sketches into full programs to allocate context across different sub-tasks, but lack mechanisms to dynamically adjust the plan based on interaction with the environment. THREAD and REPL-Plan (schroeder2025thread; liu2024interactive) instead introduce recursive substructures on top of ReAct-style agents, descending into sub-tasks to isolate local context in complex problems, but this recursion is primarily used for internal context management and the overall interaction pattern remains a standard step-by-step action loop.
It is highlighted that the core limitation across all existing paradigms remains this rigid separation of plan and action. In contrast, our work is built on the key insight that plan and action are not distinct processes, but rather represent decisions at different levels of granularity. This forms the foundation for our paradigm, which is unifying plan and action in a single representation.
@@ 4.1 Experimental Setup
We conduct experiments across three text-simulated environments that represent diverse decision-making challenges for LLM-based agents. ALFWorld (shridhar2021alfworld) provides an embodied household environment where agents navigate and manipulate objects to complete daily tasks. WebShop (yao2022webshop) simulates an online shopping scenario requiring agents to search, compare, and purchase products based on specific requirements. ScienceWorld (wang2022scienceworld) presents a scientific laboratory setting where agents conduct experiments and manipulate scientific instruments. All three environments are partially observable Markov decision processes where agents cannot fully infer the environment state from observations alone. For more detailed descriptions of the environments, please refer to Appendix A.1.
@@
For inference experiments, we compare against several mainstream paradigms: ReAct (yao2023react), which alternates between reasoning and action for iterative environment interaction; CodeAct (wang2024executable), extending ReAct’s action space from natural language to Python code. And some of the work focused on improving LLM-based agent planning: AdaPlanner (sun2023adaplanner), which pre-writes action sequences in Python and iteratively modifies them based on environmental feedback; and ADaPT (prasad2024adapt), which decomposes tasks on-demand using ReAct as the sub-task executor.
@@
For training experiments, which evaluate how well a paradigm’s data structure lends itself to learning, we conduct supervised fine-tuning (SFT) (weifinetuned) experiments on ReCode, ReAct, and CodeAct to directly compare their learning efficiency. Also, we compare some advanced training methods on the ReAct-style agent, including ETO (song2024trial), which learns from expert trajectories via Direct Preference Optimization (rafailov2023direct); and WKM (qiao2024agent), introducing parameterized world models for training on expert and synthetic trajectories. Due to differences in models, training setups, or environment configurations, we could not directly implement these latter works, but list them for reference.
@@
For inference experiments, all methods use GPT-4o mini (hurst2024gpt4o) to ensure fair comparison. We adapt few-shot prompts from prior work, making minimal modifications to fit each method’s format while keeping the number of examples consistent. The few-shot examples used for ReCode are provided in Appendix A.3.
For training experiments, we use Qwen2.5-7B-Instruct (yang2024qwen2) as the base model. We construct the training datasets through the following process: First, for each training instance, we use a powerful teacher model, DeepSeek-V3.1 (deepseekai2024deepseekv3technicalreport), to generate trajectories using ReAct, CodeAct, and ReCode respectively. Second, we filter these trajectories by keeping only the top 40% based on final reward. Third, from each retained trajectory, we extract input-output pairs as SFT training data, where the input excludes few-shot examples to focus on the model’s learning of the paradigm itself.
Table 1 presents detailed statistics of the resulting training datasets. As shown, different paradigms produce vastly different amounts of training data from the same pool of source instances. ReCode generates significantly fewer but more compact data pairs (e.g., 6,385 pairs with 4.96M tokens on ALFWorld) compared to ReAct (27,607 pairs with 18.96M tokens). All models are trained using full-parameter fine-tuning on 8 NVIDIA A800 80GB GPUs with the open-source verl framework (sheng2024hybridflow).
@@ 4.2 Inference Results
| AdaPlanner Gemini 2.5 Flash | 77.86 | 85.82 | 30.41 | 33.25 | 8.68 | 47.2 |
| ADaPT Gemini 2.5 Flash | 64.29 | 73.88 | 37.81 | 31.44 | 33.95 | 48.3 |
@@
As shown in Table 2, ReCode achieved significant performance improvements across all three environments, with an average score of 60.8, surpassing the best baseline method by approximately 10.5 (relative 20.9%). Notably, compared to the ReAct and CodeAct paradigms, ReCode improved average performance by 13.4 and 16.9, respectively. The results demonstrate ReCode’s exceptional capabilities across diverse task domains. In the ALFWorld environment, our method exhibits particularly strong performance, achieving a high score of 83.57 on seen tasks and an outstanding 96.27 on unseen tasks, substantially outperforming all baseline methods. On WebShop, ReCode maintains competitive performance with a score of 39.97, representing a 21.9% improvement over ADaPT. In the ScienceWorld environment, while ReCode shows comparable performance to ReAct on seen data, it demonstrates notable improvement on unseen tasks. These consistent improvements across varied environments underscore the effectiveness of our approach in enhancing both task-specific performance and cross-domain adaptability. We provide detailed case study in Appendix A.5, showcasing representative trajectory from ALFWorld to illustrate ReCode’s decision-making process.
@@
To illustrate that the ReCode paradigm is applicable to most models, we also experiment with Gemini Flash 2.5 (comanici2025gemini) and DeepSeek-V3.1 (deepseekai2024deepseekv3technicalreport), with results shown in Table 2. The cross-model evaluation demonstrates ReCode’s robustness and generalizability beyond a single language model architecture. On Gemini 2.5 Flash, ReCode achieved an average performance of 66.2, maintaining consistent improvements over baseline methods. Similarly, when applied to DeepSeek-V3.1, our approach obtained a 69.2 average score, further validating the paradigm’s effectiveness. These results indicate that ReCode’s performance gains are not dependent on specific model characteristics or training procedures, but rather stem from the fundamental improvements in the reasoning and code generation process.
@@
As shown in Table 3, ReCode is remarkably cost-efficient. On average, a ReCode trajectory costs 78.9% less than ReAct and 84.4% less than CodeAct. This dramatic improvement stems from ReCode’s structured exploration. Instead of the extensive, token-intensive, step-by-step reasoning cycles characteristic of reactive methods, ReCode’s hierarchical decomposition allows the agent to make more abstract and potent decisions. This leads to shorter, more direct reasoning paths, significantly reducing the total tokens required to solve a task while achieving superior performance.
@@ 4.3 Training Results
Table 5: Training data statistics and corresponding ScienceWorld performance for ReAct and ReCode. The table lists the number of data pairs and the resulting average rewards (%) for each method at different filtering percentiles ().
Data Efficiency Analysis
4.4 Evaluation on Embodied Manipulation Tasks
To further evaluate whether ReCode’s unified granularity control benefits tasks that require both high level planning and low level motor control, we conduct experiments on the ManiSkill (tao2025maniskill) benchmark. Unlike ALFWorld, WebShop, and ScienceWorld, ManiSkill requires continuous control of robot end effectors and fine grained pose adjustments, which provides a natural setting where dynamic adjustment of decision granularity is essential. We use two standard manipulation tasks, PickCube v1 and PullCube v1, based on the official ManiSkill3 Franka Panda robot. ReCode interacts with the environment through a small set of textual actions with continuous parameters; details of the wrappers, low level APIs, and task specifications are provided in Appendix A.1.
| Model | Method | PickCube | PullCube |
| GPT-4o mini | ReAct | 0 | 2 |
| CaP | 34 | 34 | |
| ReCode | 78 | 46 | |
| Gemini 2.5 Flash | ReAct | 40 | 30 |
| CaP | 82 | 44 | |
| ReCode | 90 | 100 |
We compare ReCode with two widely used paradigms for embodied decision making, ReAct and Code-as-Policies (CaP) (liang2023code), where ReAct performs step by step reasoning and emits actions, and CaP represents policies as full programs and recursively fills in missing functions before execution. We evaluate all methods with two LLMs, GPT-4o mini and Gemini 2.5 Flash, and report the success rate (%) on 50 episodes with random initialization under fixed seed for each task. The results in Table 6 show that ReAct struggles on ManiSkill, CaP improves performance but still relies on static programs that cannot adapt during execution, while ReCode consistently achieves the best results. The experiment indicate that ReCode’s unified recursive code based framework transfers to embodied manipulation and that flexible control of decision granularity yields practical benefits beyond text only environments.
@@ 6 Limitations and Future Work
While ReCode shows strong potential, its effectiveness still depends heavily on the underlying language model and the quality of the few shot examples. The framework requires the model to produce structured code in a specific format, so weaknesses in planning ability or deviations from the required format can lead to unstable performance. A natural direction for future work is to strengthen the model’s understanding of the ReCode framework through targeted pre training or fine tuning, and to explore reinforcement learning signals that reward efficient hierarchical plans and reliable expansions. It is also promising to develop mechanisms that increase robustness to generation errors and to investigate curriculum style training procedures that gradually introduce more complex tasks and reasoning patterns.